# Simplified Information Geometry Approach for Massive MIMO-OFDM Channel Estimation - Part II: Convergence Analysis

### v1

@@ II-A SIGA
and both ${\bm{\mu}}_{0}\left({\bm{\vartheta}}_{0}\right)$ and ${\bm{\Sigma}}_{0}\left({\bm{\vartheta}}_{0}\right)$ are determined by vector ${\bm{\vartheta}}_{0}$. Specifically, given ${\bm{\vartheta}}\left(t\right)=\mathbf{f}\left({\bm{\theta}}\left(t\right),{\bm{\nu}}\left(t\right)\right)$ at the $t$-th iteration, ${\bm{\vartheta}}\left(t+1\right)=\mathbf{f}\left({\bm{\theta}}\left(t+1\right),{\bm{\nu}}\left(t+1\right)\right)$ is then calculated as (II-A) [11], where $0<d\leq 1$ is the damping factor, and ${\bm{\Lambda}}\left({\bm{\nu}}\left(t\right)\right)$ and $\beta\left({\bm{\nu}}\left(t\right)\right)$ are given by
@@
Initialization: set $t=0$, set damping $d$, where $0<d\leq 1$, initialize the common NP as ${\bm{\vartheta}}\left(0\right)=\mathbf{f}\left({\bm{\theta}}\left(0\right),{\bm{\nu}}\left(0\right)\right)$ and ensure $-\frac{N-1}{\sigma_{z}^{2}}\mathbf{1}\leq{\bm{\nu}}\left(0\right)\leq\mathbf{0}$; repeat
      1. Update ${\bm{\vartheta}}=\mathbf{f}\left({\bm{\theta}},{\bm{\nu}}\right)$ as (II-A), where ${\bm{\Lambda}}\left({\bm{\nu}}\left(t\right)\right)$ and $\beta\left({\bm{\nu}}\left(t\right)\right)$ are given by (9a) and (9b), respectively;
@@ 2. $t=t+1$;
     Output: Calculate the NP of $p_{0}\left(\mathbf{h};{\bm{\vartheta}}_{0}\right)$ as ${\bm{\vartheta}}_{0}=\frac{N}{N-1}{\bm{\vartheta}}\left(t\right)$.The mean and variance of the approximate marginal, $p_{i}\left(h_{i}|\mathbf{y}\right)$, $i\in\mathcal{Z}_{M}^{+}$, are given by the $i$-th component of ${\bm{\mu}}_{0}$ and $\mathrm{diag}\left\{{\bm{\Sigma}}_{0}\right\}$, respectively, where ${\bm{\mu}}_{0}$ and ${\bm{\Sigma}}_{0}$ are calculated by (II-A).
@@ II-B Convergence
From Algorithm 1, we can find that to prove the convergence of SIGA, we only need to prove the convergence of ${\bm{\vartheta}}\left(t\right)$. Given ${\bm{\vartheta}}\left(t\right)=\mathbf{f}\left({\bm{\theta}}\left(t\right),{\bm{\nu}}\left(t\right)\right)$ at the $t$-th time, we re-express ${\bm{\vartheta}}\left(t+1\right)=\mathbf{f}\left({\bm{\theta}}\left(t+1\right),{\bm{\nu}}\left(t+1\right)\right)$ in (II-A) as the follows:
@@
where (II-B) is the iterating system of ${\bm{\nu}}$, (II-B) is the iterating system of ${\bm{\theta}}$, and the derivations are given in Appendix A. All the above matrices that need to be inverted are shown to be invertible at each iteration in Appendix A, which guarantees the iterating systems defined by (II-B) and (II-B) are valid, $\tilde{\mathbf{B}}$ and $\mathbf{B}$ are two matrix functions with ${\bm{\nu}}\left(t\right)$ being the variable, i.e., $\tilde{\mathbf{B}},\mathbf{B}:\mathbb{R}^{M}\to\mathbb{C}^{M\times M}$, and $\mathbf{b}$, $\tilde{\mathbf{g}}$ and $\mathbf{g}$ are three vector functions with ${\bm{\nu}}\left(t\right)$ being the variable, i.e., $\mathbf{b}:\mathbb{R}^{M}\to\mathbb{C}^{M}$, and $\tilde{\mathbf{g}},\mathbf{g}:\mathbb{R}^{M}\to\mathbb{R}^{M}$. Also, we have the following lemma.
@@
[…]
@@
[…]
@@
Given a finite initialization ${\bm{\theta}}\left(0\right)\in\mathbb{C}^{M\times 1}$ and ${\bm{\nu}}\left(0\right)$ with $-\frac{N-1}{\sigma_{z}^{2}}\mathbf{1}\leq{\bm{\nu}}\left(0\right)\leq\mathbf{0}$. Then, ${\bm{\theta}}\left(t\right)$ in (II-B) converges to its fixed point if the spectral radius of $\tilde{\mathbf{B}}^{\star}$ is less than $1$, i.e., $\rho\left(\tilde{\mathbf{B}}^{\star}\right)<1$, with $\rho\left(\tilde{\mathbf{B}^{\star}}\right)=\max\left\{\left|\lambda\right|:\lambda\ \textrm{is an eigenvalue of }\tilde{\mathbf{B}^{\star}}\right\}$.
@@
Given a finite initialization ${\bm{\theta}}\left(0\right)\in\mathbb{C}^{M\times 1}$ and ${\bm{\nu}}\left(0\right)$ with $-\frac{N-1}{\sigma_{z}^{2}}\mathbf{1}\leq{\bm{\nu}}\left(0\right)\leq\mathbf{0}$. Then, ${\bm{\theta}}\left(t\right)$ in (II-B) converges to its fixed point if the damping factor satisfies
@@
Given a finite initialization ${\bm{\theta}}\left(0\right)\in\mathbb{C}^{M\times 1}$ and ${\bm{\nu}}\left(0\right)$ with $-\frac{N-1}{\sigma_{z}^{2}}\mathbf{1}\leq{\bm{\nu}}\left(0\right)\leq\mathbf{0}$. Then, ${\bm{\theta}}\left(t\right)$ in (II-B) converges to its fixed point if the damping factor satisfies $d<\frac{2}{M}$.
@@ Appendix A Calculation of (II-B)
where the matrix invertibility comes from (II-A) directly.

### v3

@@ II-A SIGA
and both ${\bm{\mu}}_{0}\left({\bm{\vartheta}}_{0}\right)$ and ${\bm{\Sigma}}_{0}\left({\bm{\vartheta}}_{0}\right)$ are determined by vector ${\bm{\vartheta}}_{0}$. Specifically, given ${\bm{\vartheta}}\left(t\right)=\mathbf{f}\left({\bm{\theta}}\left(t\right),{\bm{\nu}}\left(t\right)\right)$ at the $t$-th iteration, ${\bm{\vartheta}}\left(t+1\right)=\mathbf{f}\left({\bm{\theta}}\left(t+1\right),{\bm{\nu}}\left(t+1\right)\right)$ is then calculated as (11) [11], where $0<d\leq 1$ is the damping factor, and ${\bm{\Lambda}}\left({\bm{\nu}}\left(t\right)\right)$ and $\beta\left({\bm{\nu}}\left(t\right)\right)$ are given by
@@
Initialization: set $t=0$, set damping $d$, where $0<d\leq 1$, initialize the common NP as ${\bm{\vartheta}}\left(0\right)=\mathbf{f}\left({\bm{\theta}}\left(0\right),{\bm{\nu}}\left(0\right)\right)$ and ensure $-\frac{N-1}{\sigma_{z}^{2}}\mathbf{1}\leq{\bm{\nu}}\left(0\right)\leq\mathbf{0}$;
repeat
      1. Update ${\bm{\vartheta}}=\mathbf{f}\left({\bm{\theta}},{\bm{\nu}}\right)$ as (11), where ${\bm{\Lambda}}\left({\bm{\nu}}\left(t\right)\right)$ and $\beta\left({\bm{\nu}}\left(t\right)\right)$ are given by (9a) and (9b), respectively;
@@ 2. $t=t+1$;
     Output: Calculate the NP of $p_{0}\left(\mathbf{h};{\bm{\vartheta}}_{0}\right)$ as ${\bm{\vartheta}}_{0}=\frac{N}{N-1}{\bm{\vartheta}}\left(t\right)$.The mean and variance of the approximate marginal, $p_{i}\left(h_{i}|\mathbf{y}\right)$, $i\in\mathcal{Z}_{M}^{+}$, are given by the $i$-th component of ${\bm{\mu}}_{0}$ and $\mathrm{diag}\left\{{\bm{\Sigma}}_{0}\right\}$, respectively, where ${\bm{\mu}}_{0}$ and ${\bm{\Sigma}}_{0}$ are calculated by (8).
@@ II-B Convergence
From Algorithm 1, we can find that to prove the convergence of SIGA, we only need to prove the convergence of ${\bm{\vartheta}}\left(t\right)$. Given ${\bm{\vartheta}}\left(t\right)=\mathbf{f}\left({\bm{\theta}}\left(t\right),{\bm{\nu}}\left(t\right)\right)$ at the $t$-th time, we re-express ${\bm{\vartheta}}\left(t+1\right)=\mathbf{f}\left({\bm{\theta}}\left(t+1\right),{\bm{\nu}}\left(t+1\right)\right)$ in (11) as the follows:
@@
where (12) is the iterating system of ${\bm{\nu}}$, (13) is the iterating system of ${\bm{\theta}}$, and the derivations are given in Appendix A. All the above matrices that need to be inverted are shown to be invertible at each iteration in Appendix A, which guarantees the iterating systems defined by (12) and (13) are valid, $\tilde{\mathbf{B}}$ and $\mathbf{B}$ are two matrix functions with ${\bm{\nu}}\left(t\right)$ being the variable, i.e., $\tilde{\mathbf{B}},\mathbf{B}:\mathbb{R}^{M}\to\mathbb{C}^{M\times M}$, and $\mathbf{b}$, $\tilde{\mathbf{g}}$ and $\mathbf{g}$ are three vector functions with ${\bm{\nu}}\left(t\right)$ being the variable, i.e., $\mathbf{b}:\mathbb{R}^{M}\to\mathbb{C}^{M}$, and $\tilde{\mathbf{g}},\mathbf{g}:\mathbb{R}^{M}\to\mathbb{R}^{M}$. Also, we have the following lemma.
@@
Moreover, if $\tilde{\mathbf{g}}_{\textrm{min}}\leq{\bm{\nu}}\leq\mathbf{0}$ with $\tilde{\mathbf{g}}_{\textrm{min}}\triangleq-\frac{N-1}{\sigma_{z}^{2}}\mathbf{1}\in\mathbb{R}^{M}$, we have $\tilde{\mathbf{g}}_{\textrm{min}}<\tilde{\mathbf{g}}\left({\bm{\nu}}\right)<\mathbf{0}$.
@@
Given initialization ${\bm{\nu}}\left(0\right)$ with $\tilde{\mathbf{g}}_{\textrm{min}}\leq{\bm{\nu}}\left(0\right)\leq\mathbf{0}$, where $\tilde{\mathbf{g}}_{\textrm{min}}$ is defined in Lemma 2, the sequence ${\bm{\nu}}\left(t+1\right)=\tilde{\mathbf{g}}\left({\bm{\nu}}\left(t\right)\right)$ converges to a finite fixed point ${\bm{\nu}}^{\star}$, where $\tilde{\mathbf{g}}_{\textrm{min}}<{\bm{\nu}}^{\star}<\mathbf{0}$.
@@
Given a finite initialization ${\bm{\theta}}\left(0\right)\in\mathbb{C}^{M\times 1}$ and ${\bm{\nu}}\left(0\right)$ with $-\frac{N-1}{\sigma_{z}^{2}}\mathbf{1}\leq{\bm{\nu}}\left(0\right)\leq\mathbf{0}$. Then, ${\bm{\theta}}\left(t\right)$ in (13) converges to its fixed point if the spectral radius of $\tilde{\mathbf{B}}^{\star}$ is less than $1$, i.e., $\rho\left(\tilde{\mathbf{B}}^{\star}\right)<1$, with $\rho\left(\tilde{\mathbf{B}^{\star}}\right)=\max\left\{\left|\lambda\right|:\lambda\ \textrm{is an eigenvalue of }\tilde{\mathbf{B}^{\star}}\right\}$.
@@
Given a finite initialization ${\bm{\theta}}\left(0\right)\in\mathbb{C}^{M\times 1}$ and ${\bm{\nu}}\left(0\right)$ with $-\frac{N-1}{\sigma_{z}^{2}}\mathbf{1}\leq{\bm{\nu}}\left(0\right)\leq\mathbf{0}$. Then, ${\bm{\theta}}\left(t\right)$ in (13) converges to its fixed point if the damping factor satisfies
@@
Given a finite initialization ${\bm{\theta}}\left(0\right)\in\mathbb{C}^{M\times 1}$ and ${\bm{\nu}}\left(0\right)$ with $-\frac{N-1}{\sigma_{z}^{2}}\mathbf{1}\leq{\bm{\nu}}\left(0\right)\leq\mathbf{0}$. Then, ${\bm{\theta}}\left(t\right)$ in (13) converges to its fixed point if the damping factor satisfies $d<\frac{2}{M}$.
@@ Appendix A Calculation of (12)
where the matrix invertibility comes from (9) directly.